OpenStudy (anonymous):

I GIVE MEDALS Find all the zeros of each polynomial function. f(x) = x4 + 4x3 – 6x2 – 36x – 27 Select one: a. ±1 and ±3 b. ±3 only c. ±1 only d. none of the above

OpenStudy (anonymous):

Have you tried solving this first? The first step would be to factor the polynomial, which would be \[(x-3)(x+1)(x+3)^2\]

OpenStudy (mathstudent55):

Find f(1), f(-1), f(3), f(-3)

OpenStudy (anonymous):

do you know how to go from there?

OpenStudy (anonymous):

so its alpha

OpenStudy (mathstudent55):

If f(x) = 0 for a certain x, then that x is a zero of the function.

OpenStudy (anonymous):

Yes the answer is A because when you set the binomial equal to zero, you get x = -1, x = 3, x = -3, x = -3

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

\(\large f(x) = x^4 + 4x^3 – 6x^2 – 36x – 27\) Evaluate the function for all four values of x below. For any value of x the function equals zero, that value of x is a zero of the function. \(\large f(1) = 1^4 + 4\times1^3 – 6\times1^2 – 36\times 1 – 27 =\) \(\large f(-1) = (-1)^4 + 4\times(-1)^3 – 6\times(-1)^2 – 36\times (-1) – 27 =\) \(\large f(3) = 3^4 + 4\times3^3 – 6\times3^2 – 36\times 3 – 27 =\) \(\large f(-3) = (-3)^4 + 4\times(-3)^3 – 6\times(-3)^2 – 36\times (-3) – 27 =\)

OpenStudy (mathstudent55):

@desiswag How did you factor the polynomial?